A mercury vapor lamp can be used as a 185 nm light source. What is the wavelength of the light? What is the energy of one photon in joules? Of one mole of photons in kilojoules?
Answer #1

1)
wavelength = 185 nm = 1.85*10^-7 m

2)
Find energy of 1 photon first

Given:
lambda = 1.85*10^-7 m

use:
E = h*c/lambda
=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(1.85*10^-7 m)
= 1.074*10^-18 J
Answer: 1.07*10^-18 J

3)


Energy of 1 mol = energy of 1 photon * Avogadro's number
= 1.074*10^-18*6.022*10^23 J
= 6.471*10^5 J
= 6.471*10^2 KJ
Answer: 647 KJ